# SurgicAI: A Fine-grained Platform for Data Collection and Benchmarking in Surgical Policy Learning

### v1

@@ Abstract
Despite advancements in robotic-assisted surgery, automating complex tasks like suturing remain challenging due to the need for adaptability and precision. Learning-based approaches, particularly reinforcement learning (RL) and imitation learning (IL), require realistic simulation environments for efficient data collection. However, current platforms often include only relatively simple, non-dexterous manipulations and lack the flexibility required for effective learning and generalization. We introduce SurgicAI, a novel platform for development and benchmarking addressing these challenges by providing the flexibility to accommodate both modular subtasks and more importantly task decomposition in RL-based surgical robotics. Compatible with the da Vinci Surgical System, SurgicAI offers a standardized pipeline for collecting and utilizing expert demonstrations. It supports deployment of multiple RL and IL approaches, and the training of both singular and compositional subtasks in suturing scenarios, featuring high dexterity and modularization. Meanwhile, SurgicAI sets clear metrics and benchmarks for the assessment of learned policies. We implemented and evaluated multiple RL and IL algorithms on SurgicAI. Our detailed benchmark analysis underscores SurgicAI’s potential to advance policy learning in surgical robotics. Details: https://github.com/surgical-robotics-ai/SurgicAI
@@ 1 Introduction
Robotic-assisted surgery has revolutionized the medical field by enhancing precision, reducing surgeon fatigue, and significantly shortening recovery times. However, automating complex surgical tasks such as suturing remains a significant challenge. Traditional methods [1, 2, 3, 4, 5] rely on pre-programmed instructions and heuristics, lacking the adaptability and generalization needed for diverse surgical scenarios. Additionally, there is a lack of fine-grained platforms for complex, multi-stage tasks that provide rich data and assessment metrics. These limitations highlight the need for advanced learning frameworks to handle the complexities of surgical manipulation, as well as comprehensive evaluation and benchmarking methods.
@@
In this paper, we introduce SurgicAI, which features complex multi-body dynamics, soft-body simulation and real-time interaction compatible with the da Vinci Research Kit (dVRK) [12], providing a comprehensive learning environment for developing and testing robotic suturing techniques. Moreover, SurgicAI integrates standardized pipelines, advanced hierarchical learning frameworks, diverse task suites, and benchmark performance metrics to elevate its capabilities.
The primary contributions of SurgicAI are as follows:
  • Standardized Pipeline: SurgicAI offers a comprehensive and standardized pipeline for the entire process of surgical automation. This pipeline includes collecting and preprocessing expert trajectories from teleoperation devices, training agents using reinforcement learning and imitation learning approaches, and evaluating their performance. By providing a consistent methodology, SurgicAI ensures the reliability and reproducibility of the training results.
  • Hierarchical Task Decomposition: Most platforms mentioned previously struggle with relatively complex, multi-stage tasks like suturing. SurgicAI is the first platform to implement a hierarchical learning framework specifically for managing multi-stage surgical procedures. The hierarchical structure facilitates the decomposition of intricate surgical tasks into smaller, more manageable subtasks, each handled by specialized modules. This approach not only enhances learning efficiency but also promotes the reuse of learned skills across different procedures, thereby improving the generalizability and scalability of robotic systems.
@@
  • Benchmark of Performance: SurgicAI establishes clear metrics for each suturing task and benchmarks performance across various RL and IL algorithms. With support for the Gymnasium API [13] and Stable Baseline3 (SB3) [14], it offers a user-friendly interface for defining new tasks and implementing custom algorithms. With baselines and metrics, SurgicAI also enables researchers to compare and improve their methods.
@@ 2.1 Simulation Platforms for Data Collection and Training
The development of simulation environments for data collection and training in surgical robotics has advanced significantly, driven by the need for scalable and efficient methods to train and evaluate algorithms. Frameworks such as dVRL [9] and SurRoL [10] integrate real-time physics engines with user-friendly RL libraries. Both are compatible with dVRK and have demonstrated successful policy transfer to real-world scenarios. Other works like Surgical Gym enhances data collection efficiency through high-performance GPU-based simulations, achieving significantly faster sampling rates compared to previous platforms. However, it is exclusively designed for RL implementation and does not offer interfaces for teleoperation or collecting expert demonstrations for imitation learning or demonstration-guided purposes. Additionally, most of these platforms lack soft body simulation, which limits their physical realism.
@@
Recent advancements, exemplified by ORBIT-Surgical [15], have introduced photorealistic and physics-based environments with GPU parallelization, supporting both imitation learning and reinforcement learning. Despite these improvements, such platforms are limited to relatively simple tasks like reaching and placing, and they struggle with more complex, multi-stage tasks such as suturing. In contrast, SurgicAI offers a comprehensive solution by addressing these limitations and providing robust support for complex suturing procedures.
2.2 Learning Based Approach in Surgical Automation
@@
Robot learning has been applied to various surgical scenarios, including debris removal [16, 17, 18], tissue retraction [19, 20, 21], suturing [22, 23, 24], and cutting [25, 26]. The results from these applications show promising potential for machine learning-based approaches to become key solutions for surgical automation. However, traditional RL methods face significant challenges in these contexts due to sparse rewards, discounting issues, and exploration difficulties inherent in high-dimensional and long-horizon tasks [27, 28]. These challenges often hinder RL algorithms from effectively learning and performing complex surgical tasks. To address these issues, many existing methods rely heavily on elaborate reward shaping to facilitate extensive exploration [19, 29]. This process typically involves subjective manual engineering, which can introduce biases and lead to unexpected policies that get stuck in local optima [28]. The reliance on handcrafted rewards can also limit the generalizability of the learned policies, as they may not perform well in slightly different or unforeseen scenarios.
@@
Advanced methodologies, such as [30, 31], have been developed to enhance exploration efficiency by incorporating demonstration data into RL. This approach allows learning from knowledge from experts, reducing the time and computational resources needed to achieve effective policy learning. Works from [32, 33, 24] predefine primitive skills and sequence them at transition points, creating a more structured and guided learning process. This method helps break down complex tasks into simpler, more manageable components, which can be learned more effectively.
@@
Building on these advancements, the hierarchical framework [34, 35, 36], utilized in SurgicAI, has been introduced to better coordinate sub-policies. This hierarchical approach divides the overall task into a hierarchy of sub-tasks, each with its own specific policies. By structuring the learning process in this way, the hierarchical framework not only improves learning efficiency but also enhances the robustness and adaptability of the system. Moreover, SurgicAI distinctively pioneers the application of image-guided imitation learning algorithms for complex manipulative tasks involving patient-side manipulators (PSMs). While platforms such as [10] and [11] have defined some basic ECM tracking tasks using image data, and [15] demonstrates the capability for collecting RGB or segmented image data, they have not extensively applied imitation learning to the intricate and precise manipulations required for tasks like suturing.
Figure 1: Detailed Workflow of our Simulation Environment. The ADF files define various simulation objects, such as rigid bodies, joints, and cameras. The launch file specifies simulation parameters and model-specific details. The AMBF Simulator processes these parameters to initialize the environment, while the AMBF Client bridges the simulator and the control scripts via ROS topics. Method APIs provide essential functions for controlling the simulation, including kinematics and servo control, while teleoperation scripts enable connections with multiple devices. The Gymnasium API integrates RL and IL capabilities, optimizing policies through interaction with the control scripts.
@@ 3.1 Simulation Environment
The simulation environment of SurgicAI is based on the AccelNet Surgical Robotics Challenge (SRC) [37] and utilizes the Asynchronous Multi-Body Framework (AMBF) [38]. This environment includes two Patient Side Manipulators (PSMs) from the da Vinci Surgical System, an Endoscopic Camera Manipulator (ECM), a needle with or without thread, and realistic phantoms marked with entry and exit holes for suturing. A key feature of this environment is its support for soft body dynamics, which significantly enhances the realism of tissue simulations. The system is fully compatible with the dVRK system, enabling real-time communication with various teleoperation devices such as dVRK MTMs, Geomagic, and Razer Hydra. Additionally, our environment supports a headless mode to facilitate efficient data sampling without the computational overhead of graphical rendering. The system generates multi-modal data, including ground truth positions, RGB images, and labeled depth maps, making it an excellent platform for data-driven approaches to suturing automation. The simulator workflow is detailed in Fig. 1.
@@ 3.2 Environment Settings
Our simulation environment adheres to the Gymnasium framework rules, ensuring compatibility with various RL and IL tasks.
Observation and State Space: This environment features a versatile observation space that accommodates both low-dimensional and high-dimensional states for comprehensive data collection. In low-dimensional settings, the observation space includes ground-truth 6D poses of objects such as needles, entry and exit holes, and end-effectors. When using Hindsight Experience Replay (HER) [39], the agent receives extra lists of achieved and desired goals, facilitating learning from past experiences. High-dimensional settings, on the other hand, comprise scene-invariant RGB images from multiple views and proprioception data from PSMs, specifically the 6D pose of the end-effector in its local frame. Our experiments utilize low-dimensional states in RL methods for quick verification, whereas high-dimensional states are employed only for imitation learning, leveraging detailed visual and proprioceptive information.
@@ 3.3 Task Overview
To address the complexity of the suturing process, we have segmented the process into several distinct subtasks, similar to [33]: approaching and grasping the needle, placing the needle at the entry hole, inserting the needle, handing off the needle (regrasping it with the other PSM), and pulling out the needle. This segmentation is informed by empirical observations that show users tend to slow down during transitions between these tasks, making it easier to define and separate each subtask. For specific low-dimensional settings, each subtask is defined as below; while the settings described below offer functionally plausible thresholds for successful task execution, the numerical thresholds for distances and orientation errors are also flexible to adapt to different levels of precision.
Grasping: At the beginning of each episode, the needle’s position is randomized. A specific point on the needle is designated as the grasping point. The episode is considered successful if the remaining distance between the jaw and the grasping point is within 1 mm, the orientation error is within 10 degrees, and the ray-tracing sensor attached to the gripper successfully detects the needle object.
Placing: The initial state involves the jaw randomly grasping the needle with with positional and orientation errors of up to 1.5 mm and 12 degrees respectively from the expected grasping orientation. Eventually, the needle tip should align with the entry point’s center and perpendicular to the entry plane. Success criteria are a distance within 5 mm and an orientation error within 10 degrees.
@@
Handoff: Starting with the needle passing through the phantom, a grasping point on the needle is identified. Success is achieved if the remaining distance between the jaw and the grasping point is within 2mm, the orientation error is within 15 degrees, and the ray-tracing sensor detects the needle.
Pullout: The initial state involves the jaw randomly grasping the needle with a positional error up to 1.5mm and an orientation deviation up to 20 degrees from the expected grasping point in the handoff task. The goal is to fully extract the needle from the phantom and reach a predetermined end point. Success criteria are a distance within 5mm and an orientation error within 20 degrees.
@@ 3.4 Hierarchical Architecture for Multi-stage Tasks
High-Level Policy (HLP) At the core of the hierarchical architecture is the High-Level Policy. The HLP is responsible for sequencing and coordinating the Low-Level Policy. Based on the current state of the task, the HLP determines the appropriate subtask to activate, managing transitions between subtasks and ensuring that each stage of the task is executed in the correct order. This component allows the system to adapt to dynamic changes in the environment and varying initial configurations.
Low-Level Policy (LLP) The Low-Level Policies are specialized neural networks trained to execute specific subtasks. Each LLP is responsible for a particular aspect of the overall task, such as Grasping, Placing, Inserting, Handoff, or Pullout. These subtasks are parameterized and trained independently, enabling the LLPs to master their respective tasks efficiently. The specialization of LLPs enables focused learning and optimization, achieving more robust and reliable performance for each subtask.
Integration and Coordination The hierarchical framework operates in a cyclical manner to ensure efficient task execution: 1) The HLP generates a subtask based on the current state of the environment; 2) The corresponding LLP executes the subtask and reaches its terminal state, upon which the process loops back to the HLP for the next subtask generation. This process continues iteratively until the overall task is completed. By breaking down complex tasks into simpler components, the hierarchical architecture ensures that each subtask is handled by a network specialized in that particular operation. The HLP’s role in managing the sequence and coordination of subtasks is crucial for maintaining the overall coherence and efficiency of the task execution.
@@ 4.1 Platform Usage Guideline
Figure 3 illustrates the pipeline for RL agents. The pipeline consists of three main stages: Initialization, Training, and Evaluation. The initialization stage begins with setting up ROS and SRC, initializing the Gymnasium environment which defines reset, step, and reward functions for each task, and defining learning algorithms. Additionally, it involves configuring the random seed, maximum episode length, network architecture, replay buffer, and other hyper-parameters. The framework can be adapted for imitation learning by adjusting the parameters accordingly.

### v3

# SurgicAI: A Hierarchical Platform for Fine-Grained Surgical Policy Learning and Benchmarking

@@ Abstract
Despite advancements in robotic-assisted surgery, automating complex tasks like suturing remains challenging due to the need for adaptability and precision. Learning-based approaches, particularly reinforcement learning (RL) and imitation learning (IL), require realistic simulation environments for efficient data collection. However, current platforms often include only relatively simple, non-dexterous manipulations and lack the flexibility required for effective learning and generalization. We introduce SurgicAI, a novel platform for development and benchmarking that addresses these challenges by providing the flexibility to accommodate both modular subtasks and more importantly task decomposition in RL-based surgical robotics. Compatible with the da Vinci Surgical System, SurgicAI offers a standardized pipeline for collecting and utilizing expert demonstrations. It supports the deployment of multiple RL and IL approaches, and the training of both singular and compositional subtasks in suturing scenarios, featuring high dexterity and modularization. Meanwhile, SurgicAI sets clear metrics and benchmarks for the assessment of learned policies. We implemented and evaluated multiple RL and IL algorithms on SurgicAI. Our detailed benchmark analysis underscores SurgicAI’s potential to advance policy learning in surgical robotics. Details: https://github.com/surgical-robotics-ai/SurgicAI
@@ 1 Introduction
Robotic-assisted surgery (RAS) has revolutionized the medical field by enhancing precision, reducing surgeon fatigue, and significantly shortening recovery times. While existing RAS systems lack automation, arguably automation has the potential to improve upon the current standard of care and surgeon workload. However, automating complex surgical tasks such as suturing remains a significant challenge. Traditional methods [1, 2, 3, 4, 5] rely on pre-programmed instructions and heuristics, lacking the adaptability and generality needed for diverse surgical scenarios. Additionally, there is a lack of fine-grained platforms for complex, multi-stage tasks that provide rich data and assessment metrics. These limitations highlight the need for advanced learning frameworks to handle the complexities of surgical manipulation, as well as comprehensive evaluation and benchmarking methods.
@@
In this paper, we introduce SurgicAI, which features complex deformable thread simulation, and real-time interaction compatible with the da Vinci Research Kit (dVRK) [12], providing a comprehensive learning environment for developing and testing robotic suturing techniques. Moreover, SurgicAI integrates standardized pipelines, advanced hierarchical learning frameworks, diverse task suites, and benchmark performance metrics to evaluate its capabilities.
The primary contributions of SurgicAI are the following:
  • Standardized Pipeline: SurgicAI offers a comprehensive and standardized pipeline for the entire process of surgical automation. This pipeline includes collecting and preprocessing expert trajectories from teleoperation devices, training agents using reinforcement learning and imitation learning approaches, and evaluating their performance.
  • Hierarchical Task Decomposition: Most platforms mentioned previously [9, 10, 11] struggle with relatively complex, multi-stage tasks like suturing. SurgicAI is the first platform to implement an open-source hierarchical learning framework specifically for managing multi-stage surgical procedures. The hierarchical structure facilitates the decomposition of intricate surgical tasks into smaller, more manageable subtasks, each handled by specialized modules. This approach not only enhances learning efficiency but also promotes the reuse of learned skills across different procedures, thereby improving the generality and scalability of robotic systems. Figure 1 shows how we can train both high-level policy and low-level policies to conduct the multi-stage procedures.
@@
  • Benchmark of Performance: SurgicAI establishes clear metrics for each suturing task and benchmarks performance across various RL and IL algorithms. With support for the Gymnasium API [13], and RL libraries like Stable Baseline3 (SB3) [14] and d3rlpy [15], SurgicAI offers a user-friendly interface for defining new tasks and implementing custom algorithms. With baselines and metrics, SurgicAI also enables researchers to compare and improve their methods.
@@ 2.1 Simulation Platforms for Data Collection and Training
The development of simulation environments for data collection and training in surgical robotics has advanced significantly, driven by the need for scalable and efficient methods to train and evaluate algorithms. Frameworks such as dVRL [9] and SurRoL [10] integrate real-time physics engines with user-friendly RL libraries. Both are compatible with dVRK and have demonstrated successful policy transfer to real-world scenarios. Other work like Surgical Gym enhances data collection efficiency through high-performance GPU-based simulations, achieving significantly faster sampling rates compared to previous platforms. However, it is exclusively designed for RL implementation and does not offer interfaces for teleoperation or collecting expert demonstrations for imitation learning or demonstration-guided purposes. Additionally, most simulations lack soft body simulation, except for recent developments in SurRoL [16]. This limits their physical realism.
@@
Recent advancements, exemplified by ORBIT-Surgical [17], have introduced photorealistic and physics-based environments with GPU parallelization, supporting both imitation learning and reinforcement learning. Despite these improvements, such platforms are limited to relatively simple tasks like reaching and placing, and they struggle with more complex, multi-stage tasks such as suturing. In contrast, the AccelNet Surgical Robotics Challenge (SRC) [18] offers a comprehensive solution by addressing these limitations and providing robust support for complex suturing procedures. Our work provides the necessary interface and library support to train and test RL and IL algorithms with low-level dynamics provided in SRC, using a hierarchical framework for long-term planning.
2.2 Learning Based Approaches in Surgical Automation
@@
Robot learning has been applied to various surgical scenarios, including debris removal [19, 20, 21], tissue retraction [22, 23, 24], suturing [25, 26, 27], and cutting [28, 29]. The results from these applications show promising potential for machine learning-based approaches to become key solutions for surgical automation. However, traditional RL methods face significant challenges in these contexts due to sparse rewards, discounting issues, and exploration difficulties inherent in high-dimensional and long-horizon tasks [30, 31]. These challenges often hinder RL algorithms from effectively learning and performing complex surgical tasks. To address these issues, many existing methods rely heavily on elaborate reward shaping to facilitate extensive exploration [22, 32]. This process typically involves subjective manual engineering, which can introduce biases and lead to unexpected policies that get stuck in local optima [31]. The reliance on handcrafted rewards can also limit the generality of the learned policies, as they may not perform well in slightly different or unforeseen scenarios.
@@
Advanced methodologies, such as [33, 34], have been developed to enhance exploration efficiency by incorporating demonstration data into RL. This approach allows learning from expert knowledge, reducing the time and computational resources needed to achieve effective policy learning. Work from [35, 36, 27] predefined primitive skills and sequences them at transition points, creating a more structured and guided learning process. These methods help break down complex tasks into simpler, more manageable components, which can be learned more effectively.
@@
Building on these advancements, the hierarchical framework [37, 38, 39], utilized in SurgicAI, has been introduced to better coordinate sub-policies. This hierarchical approach divides the overall task into a hierarchy of sub-tasks, each with its own specific policies. By structuring the learning process in this way, the hierarchical framework not only improves learning efficiency but also enhances the robustness and adaptability of the system. Moreover, SurgicAI distinctively pioneers the application of image-guided imitation learning algorithms for complex manipulative tasks involving patient-side manipulators (PSMs). While platforms such as [10] and [11] have defined some basic Endoscopic Camera Manipulator (ECM) tracking tasks using image data, and [17] demonstrated the capability for collecting RGB or segmented image data, they have not extensively applied imitation learning to the intricate and precise manipulations required for tasks like suturing.
@@ 3.1 Simulation Environment
The simulation environment of SurgicAI is based on the SRC [18] and utilizes the Asynchronous Multi-Body Framework (AMBF) [40, 41]. This environment includes two Patient Side Manipulators (PSMs) from the da Vinci Surgical System, an Endoscopic Camera Manipulator (ECM), a needle with or without thread, and realistic phantoms marked with entry and exit holes for suturing. A key feature of this environment is the simulation of a deformable thread, its interaction with the phantom, and the ability to pass it through the specified holes in the phantom, which significantly enhances the realism of tissue simulations. The system is fully compatible with the dVRK system, enabling real-time communication with various teleoperation devices such as dVRK MTMs, Geomagic, and Razer Hydra. Additionally, our environment supports a headless mode to facilitate efficient data sampling without the computational overhead of graphical rendering. The system generates multi-modal data, including ground truth positions, RGB images, and labeled depth maps, making it a suitable platform for data-driven approaches to suturing automation. The workflow is detailed in Figure 2.
Figure 2: Detailed Workflow of our Simulation Environment. The AMBF Description Files (ADF) define various simulation objects, such as rigid bodies, joints, and cameras. The launch file specifies simulation parameters and model-specific details. The AMBF Simulator processes these parameters to initialize the environment, while the AMBF Client bridges the simulator and the control scripts via ROS topics. Method APIs provide essential functions for controlling the simulation, including kinematics and servo control, while teleoperation scripts enable connections with multiple devices. The Gymnasium API integrates RL and IL capabilities, optimizing policies through interaction with the control scripts.
@@ 3.2 Environment Settings
Our simulation environment adheres to the Gymnasium framework rules [13], ensuring compatibility with various RL and IL tasks.
Observation and State Space: This environment features a versatile observation space that accommodates both low-dimensional and high-dimensional states for comprehensive data collection. In low-dimensional settings, the observation space includes ground-truth 6D poses of objects such as needles, entry and exit holes, and end-effectors. When using Hindsight Experience Replay (HER) [42], the agent receives extra lists of achieved and desired goals, facilitating learning from past experiences. High-dimensional settings, on the other hand, comprise scene-invariant RGB images from multiple views and proprioception data from PSMs, specifically the 6D pose of the end-effector in its local frame. Our settings support the online and offline training for both low and high dimensional state space. In our experiments, we utilize low-dimensional states in online and offline RL for rapid validation, whereas high-dimensional states are reserved for offline imitation learning, as it allows us to bypass the time-intensive process of collecting online image data.
@@ 3.3 Task Overview
To address the complexity of the suturing process, we have segmented the process into several distinct subtasks, similar to [36]: approaching and grasping the needle, placing the needle at the entry hole, inserting the needle, handing off the needle (regrasping it with the other PSM), and pulling out the needle. This segmentation is informed by empirical observations that show users tend to slow down during transitions between these tasks, making it easier to define and separate each subtask. For specific low-dimensional settings, each subtask is defined as below. While the settings described below offer functionally plausible thresholds for successful task execution, the numerical thresholds for distances and orientation errors are also flexible to adapt to different levels of precision.
Grasping: At the beginning of each episode, the needle’s position is randomized. A specific point on the needle is designated as the grasping point. The episode is considered successful if the remaining distance between the jaw and the grasping point is within 1 mm, the orientation error is within 10 degrees, and the contact sensor attached to the gripper successfully detects the needle object.
Placing: The initial state involves the jaw randomly grasping the needle with positional and orientation errors of up to 1.5 mm and 12 degrees respectively from the expected grasping orientation. Eventually, the needle tip should align with the entry point’s center and perpendicular to the entry plane. Success criteria are a distance within 5 mm and an orientation error within 10 degrees.
@@
Handoff: Starting with the needle passing through the phantom, a grasping point on the needle is identified. Success is achieved if the remaining distance between the jaw and the grasping point is within 2 mm, the orientation error is within 15 degrees, and the contact sensor detects the needle.
Pullout: The initial state involves the jaw randomly grasping the needle with a positional error of up to 1.5 mm and an orientation deviation up to 20 degrees from the expected grasping point in the handoff task. The goal is to fully extract the needle from the phantom and reach a predetermined end point. Success criteria are a distance within 5 mm and an orientation error within 20 degrees.
@@ 3.4 Hierarchical Architecture for Multi-stage Tasks
High-Level Policy (HLP): At the core of the hierarchical architecture is the High-Level Policy, which utilizes a neural network for decision making. The HLP is responsible for sequencing and coordinating the Low-Level Policy. Based on the current state of the task, the HLP determines the appropriate subtask to activate, managing transitions between subtasks and ensuring that each stage of the task is executed in the correct order. This component allows the system to adapt to dynamic changes in the environment and varying initial configurations.
Low-Level Policy (LLP): The Low-Level Policies are specialized neural networks trained to execute specific subtasks. Each LLP is responsible for a particular aspect of the overall task, such as grasping, placing, inserting, handoff, or pullout. These subtasks are parameterized and trained independently, enabling the LLPs to master their respective tasks efficiently. The specialization of LLPs enables focused learning and optimization, achieving more robust and reliable performance for each subtask.
Integration and Coordination: The hierarchical framework operates in a cyclical manner to ensure efficient task execution: 1) The HLP generates a subtask based on the current state of the environment; 2) The corresponding LLP executes the subtask and reaches its terminal state, upon which the process loops back to the HLP for the next subtask generation. This process continues iteratively until the overall task is completed. By breaking down complex tasks into simpler components, the hierarchical architecture ensures that each subtask is handled by a network specialized in that particular operation. The HLP’s role in managing the sequence and coordination of subtasks is crucial for maintaining the overall coherence and efficiency of the task execution.
@@ 4.1 Platform Usage Guideline
Figure 3 illustrates the training pipeline for RL agents. The pipeline consists of three main stages: Initialization, Training, and Evaluation. The initialization stage begins with setting up ROS and SRC, initializing the Gymnasium environment which defines reset, step, and reward functions for each task, and defining learning algorithms. Additionally, it involves configuring the random seed, maximum episode length, network architecture, replay buffer, and other hyper-parameters. The framework can be adapted for imitation learning by adjusting the parameters accordingly.